# Towards Backdoor Stealthiness in Model Parameter Space

### v2

@@ Abstract.
Model Inspection NC (Wang et al., 2019)
Tabor (Guo et al., 2020)
FeatureRE (Wang et al., 2022a)
Unicorn (Wang et al., 2023)
BTI-DBF (Xu et al., 2024a)
Input Inspection Scale-up (Guo et al., 2023)
IBD-PSC (Hou et al., 2024)
CT (Qi et al., 2023b)
Pruning FP (Liu et al., 2018a)
ANP (Wu and Wang, 2021)
CLP (Zheng et al., 2022)
RNP (Li et al., 2023a)
@@ 1. Introduction
While deep neural networks (DNNs) have achieved excellent performance on various tasks, they are vulnerable to backdoor attacks. Backdoor attacks insert a secret functionality into a model, which is activated by malicious inputs during inference. Such inputs contain an attacker-chosen property called the trigger. Backdoored DNNs can be created by training with poisoned data (Gu et al., 2019; Chen et al., 2017; Nguyen and Tran, 2021; Turner et al., 2019). More powerful and stealthy backdoors can also be injected through the control of a training process (Shokri et al., 2020; Bagdasaryan and Shmatikov, 2021; Cheng et al., 2021; Zhong et al., 2022; Qi et al., 2022; Mo et al., 2024; Zhao et al., 2022), or by direct weights modification of the victim model (Hong et al., 2022; Cao et al., 2024).
@@
In early backdoor attacks (Gu et al., 2019; Chen et al., 2017; Liu et al., 2018b), triggers could induce noticeable changes that human inspectors or anomaly detectors (Chen et al., 2018; Wang et al., 2019; Liu et al., 2019) could easily spot. To enhance the ability to remain undetected against such defenses (i.e., achieve input-space stealthiness), smaller or more semantic-aware triggers are designed (Nguyen and Tran, 2021; Doan et al., 2021b; Wang et al., 2022b). Input-space stealthy backdoor attacks usually need to change labels of poisoned samples to the target class (i.e., dirty-label), which makes detection easier (Chen et al., 2018). To this end, another line of backdoor attacks poisons the training data without changing the labels (Turner et al., 2019; Zeng et al., 2023) (i.e., clean-label), improving backdoor stealthiness.
@@
Despite the stealthiness concerning input images and labels, it has been widely observed that existing backdoor attacks introduce separable representations in the feature space, which can be exploited to develop backdoor defenses (Wang et al., 2022a; Qi et al., 2023a; Min et al., 2023; Zhu et al., 2023; Xu et al., 2024a). For example, featureRE (Wang et al., 2022a) utilizes feature separability and designs a feature space constraint to reverse-engineer the backdoor trigger. In response to feature-space defenses, state-of-the-art (SOTA) backdoor attacks focus on eliminating the separability in the feature space (Mo et al., 2024; Qi et al., 2023a; Zhu et al., 2024; Shokri et al., 2020) to increase the feature-space stealthiness, i.e., the undetectability against feature-space defenses. Considering a different threat model, supply-chain backdoor attacks assume control over the training or directly modify the model’s weights (Cheng et al., 2021; Hong et al., 2022; Cao et al., 2024), and the backdoored model is provided as a service or as the final product. For example, supply-chain attacks could introduce a penalty to the training loss that decreases the distance between the backdoor and benign features to increase feature-space stealthiness (Shokri et al., 2020; Zhao et al., 2022; Doan et al., 2021a; Zhong et al., 2022).
@@
An important observation is that most backdoor attacks are designed to be stealthy to resist a specific type of defense. For example, WaNet (Nguyen and Tran, 2021) and Bpp (Wang et al., 2022b) design imperceptible triggers to bypass input-space defenses (such as NC (Wang et al., 2019)), but introduce significant separability in the feature space (Wang et al., 2022a). Adap-patch (Qi et al., 2023a) avoids feature separability but uses patch-based triggers, which a human inspector can detect. More critically, current backdoor attacks are barely evaluated against parameter-space defenses (Zhu et al., 2023; Lin et al., 2024; Xu et al., 2024b; Wu and Wang, 2021; Zheng et al., 2022; Li et al., 2023a). This oversight is significant because backdoor behaviors are ultimately embedded in and reflected by the parameters of the backdoored model, which is the final product of any backdoor attack. As such, there is a lack of systematic evaluation of backdoor attacks against the latest parameter-space defenses.
@@
Figure 1. Diagram illustrating the working mechanism of Grond. On the left, universal PGD (UPGD) perturbation is generated as backdoor patterns to be injected. In the middle, ABI is applied where perturbed samples are iteratively used to train the model, and the model parameters are pruned to limit the magnitude of prominent backdoored weights. On the right, the output backdoored model that considers comprehensive stealthiness is deployed, where 1) the triggers are invisible, 2) the features of trigger samples are inseparable, and 3) the backdoored model weights are hardly distinguishable from benign model weights. Perturbations generated by UPGD are scaled up 10×10$\times$10 × for visualization.
@@ 2.1. Preliminaries on Backdoor Training
In the inference stage, the trained $f$ performs well on benign data $\hat{{\bm{x}}}$, but predicts $G_{x}(\hat{{\bm{x}}})$ as $G_{y}(\hat{y})$.
@@ 2.2. Backdoor Attacks
Backdoor attacks compromise the integrity of the victim model so that the model performs naturally on benign inputs but is misled to the target class by inputs containing the backdoor trigger. The trigger can be a visible pattern inserted into the model’s input in the input space or a property that affects the feature representation of the model’s input in the feature space. Eventually, however, the backdoored model’s parameters in the parameter space will be altered regardless of the exact backdoor attack (see Figure 2). To insert a backdoor, the attacker is assumed to only control a small portion of the training data under the poison training scenario (Gu et al., 2019; Chen et al., 2017; Zhang et al., 2021). In the supply-chain setting (backdoor models provided to users), the attacker also controls the training process (Shokri et al., 2020; Nguyen and Tran, 2020; Bagdasaryan and Shmatikov, 2021; Nguyen and Tran, 2021; Wang et al., 2022b). Moreover, the backdoor can also be created by directly modifying the model’s weights (Liu et al., 2017; Hong et al., 2022; Qi et al., 2022; Cao et al., 2024).
@@
Input-space attacks. Traditional attacks typically use simple patterns as their triggers. For example, BadNets (Gu et al., 2019) uses a fixed patch, and Blend (Chen et al., 2017) mixes a Hello Kitty pattern into the images as the trigger. These non-stealthy triggers introduce abnormal data into training data and can be easily detected by human inspectors or defenses (Chen et al., 2018; Wang et al., 2019). To improve the stealthiness, various triggers are proposed to achieve invisibility in the input space. IAD (Nguyen and Tran, 2020) designed a dynamic solution in which the triggers vary among different inputs. WaNet (Nguyen and Tran, 2021) proposed the warping-based trigger, which is invisible to human inspection. Although these methods successfully build invisible triggers and bypass traditional defenses (Wang et al., 2019), they still introduce separable features and can be detected by feature-space defenses (Wang et al., 2022a; Xu et al., 2024a). These input-invisible attacks can be even more noticeable than input-visible attacks (BadNet, Blend) in the feature space (Xu et al., 2024b). We conjecture this is because they have fewer modifications on input pixels than input-visible attacks. Therefore, input-invisible attacks require more influential features to achieve a successful attack.
@@
Feature-space attacks. Knowing the vulnerability of input-space attacks against feature-space defenses, backdoor attacks are improved for feature-space stealthiness. A common threat model of this attack type is to assume additional control over the training process. For example, (Shokri et al., 2020; Doan et al., 2021a; Zhao et al., 2022; Zhong et al., 2022) directly designed loss functions to minimize the difference between the backdoor and benign features. Aside from design loss penalties, TACT (Tang et al., 2021) and SSDT (Mo et al., 2024) point out that source-specific (poison only the specified source classes) attack helps to obscure the difference in features between benign and backdoor samples. In addition, (Qi et al., 2023a) proposed Adap-blend and Adap-patch, which obscures benign and backdoor features by 1) including poisoned samples with the correct label, 2) asymmetric triggers (using a stronger trigger at inference time), and 3) trigger diversification (using diverse variants of the trigger during training). Unfortunately, existing attacks lack systematic evaluation against the latest defenses. For example, Adap-blend can be thoroughly mitigated by recent works (Zhu et al., 2023; Xu et al., 2024a, b). In summary, feature-space attacks usually introduce visible triggers and cannot defeat the latest defenses.
@@
Supply-chain attacks. Supply-chain attacks are getting more attention due to their potential in real-world applications where backdoored models are provided as the final product to users. In supply-chain attacks, adversaries could control both training data and the training process. Note that feature-space attacks (Shokri et al., 2020; Lin et al., 2020; Doan et al., 2021a; Zhao et al., 2022; Ren et al., 2021; Cheng et al., 2021; Zhong et al., 2022; Xia et al., 2023; Mo et al., 2024) with the assumption of control over the training process are a subset of supply-chain attacks, as their output is the backdoor model. In addition to training control, another kind of supply-chain attack directly adjusts the model’s weights in parameter space to introduce a backdoor, i.e., parameter-space attack. T-BFA (Rakin et al., 2022), TBT (Rakin et al., 2020), and ProFlip (Chen et al., 2021) explore modifying a sequence of susceptible bits of DNN parameters stored in the main memory (e.g., DRAM) to inject the backdoor. SRA (Qi et al., 2022) and handcrafted Backdoor (Hong et al., 2022) directly modify a subset of models’ parameters to increase the logits of the target class. However, these attacks require a local benign dataset to guide the search for the subset of parameters to be modified. Data-free backdoor (Lv et al., 2023) releases the requirement of benign data by collecting substitute data irrelevant to the main task and fine-tuning using the substitute data. DFBA (Cao et al., 2024) further proposes a retraining-free and data-free backdoor attack by injecting a backdoor path (a single neuron from each layer except the output layer) into the victim model. In summary, supply-chain attacks focus on increasing the backdoor’s effectiveness without comprehensively considering parameter-space defenses.
@@ 2.3. Backdoor Defenses
Backdoor defenses can be classified into detection and mitigation. Detection refers to determining whether a model is backdoored (model detection(Wang et al., 2019; Liu et al., 2019; Zhao et al., 2022; Wang et al., 2023; Xu et al., 2024a) or a given input is applied with a trigger (input detection(Gao et al., 2019; Guo et al., 2023; Mo et al., 2024). Mitigation refers to erasing the backdoor effect from the victim model by pruning the backdoor-related neurons (pruning-based defenses) (Liu et al., 2018a; Wu and Wang, 2021; Zheng et al., 2022; Li et al., 2023a) or unlearning the backdoor trigger (fine-tuning-based defenses) (Zhu et al., 2023; Zeng et al., 2022; Min et al., 2023; Xu et al., 2024a). In addition, recent works (Li et al., 2021; Qi et al., 2023b; Wei et al., 2024) also consider the home-field advantage [^1] The defender has full control of the system and could access the training process. to design more powerful proactive defenses.
@@
Backdoor detection. Backdoor trigger reverse engineering (also known as trigger inversion) is considered one of the most practical defenses for backdoor detection as it can be applied to both poisoning training and supply-chain scenarios (Wang et al., 2022a, 2023; Xu et al., 2024a, b), i.e., it is a post-training method. Specifically, trigger inversion works by searching for a potential backdoor trigger for a specific model. The model is determined as backdoored if a trigger is found, and the trigger can be used to unlearn the backdoor. The searching is implemented as an optimization process corresponding to the model and a local benign dataset. For example, NC (Wang et al., 2019) firstly proposes trigger inversion for detection by optimizing the mask and pattern in the input space that can mislead the victim to the target class. This optimization is repeated for all classes. The model is considered backdoored if an outlier significantly smaller than the triggers for all other classes exists. Although methods similar to NC perform well against fixed patch trigger attacks, such as BadNets (Gu et al., 2019) and Blend (Chen et al., 2017), they may not be effective against input-stealthy attacks like WaNet (Nguyen and Tran, 2021). To address this problem, FeatureRE (Wang et al., 2022a) moves trigger inversion from input space to feature space. Unicorn (Wang et al., 2023) further proposes a transformation function for attacks in other spaces, such as numerical space (Wang et al., 2022b). Recent works (Xu et al., 2024a, b) focus on exploring new optimization objectives that address the inefficiency problem of previous trigger inversion methods due to optimization over all classes. BTI-DBF (Xu et al., 2024a) trains a trigger generator by maximizing the backdoor feature difference between benign samples and their generated version (by the trigger generator) and minimizing the benign feature difference. BAN (Xu et al., 2024b) optimizes the noise on neuron weights rather than input pixels to activate the potential backdoor, which further improves both effectiveness and efficiency.
@@
Backdoor mitigation. Backdoor mitigation consists of fine-tuning and pruning, which are effective and do not assume knowledge of backdoor triggers. Pruning methods aim to find and remove backdoor-related neurons. FP (Liu et al., 2018a) eliminates dormant neurons on benign inputs and then fine-tunes the pruned network. ANP (Wu and Wang, 2021) searches for backdoor-related neurons by adding adversarial noise to neuron weights to activate the backdoor. RNP (Li et al., 2023a) uses an unlearning and recovering process on benign data to expose backdoor neurons, as the recovering will force the backdoor neurons to be silent for the main benign task. Unlike these pruning methods guided by benign data, CLP (Zheng et al., 2022) directly analyzes the Channel Lipschitzness Constant of the network and prunes the high Lipschitz constant channels in a data-free manner.
@@
Traditional fine-tuning as a defense usually needs trigger inversion methods to recover the trigger and then unlearn the trigger. For example, BTI-DBF(U) (Xu et al., 2024a) fine-tunes backdoor models using triggers recovered by their inversion algorithm. However, there is no guarantee that the recovered trigger is the true trigger for the backdoor. Recent works also consider fine-tuning without the trigger information but with prior human knowledge. For example, FT-SAM (Zhu et al., 2023) observes a positive correlation between the weight norm of neurons and backdoor-related neurons. Then, they propose a fine-tuning method to revise the large outliers of weight norms using Sharpness-Aware Minimization (SAM). I-BAU (Zeng et al., 2022) forms a min-max fine-tuning similar to adversarial training, where the inner maximizing searches for perturbations that mislead the model, and the outer minimizing is to keep the model’s capability on benign data. FST (Min et al., 2023) assumes the backdoor and benign features should be disentangled and actively shifting features while fine-tuning by encouraging the discrepancy between the original backdoor model and the fine-tuned model.
@@
Proactive defense. Several methods have been proposed to exploit the home-field advantage, i.e., a stronger defender, for better defensive performance. ABL (Li et al., 2021) proposes two techniques to avoid learning the backdoor task while training on the poisoned data: 1) trapping the loss value of each example around a certain threshold because backdoor tasks are learned much faster than the main task, and their loss decreases much faster. The samples with lower loss are recorded as poisoned samples; 2) unlearning the backdoor with the recorded poisoned samples. CT (Qi et al., 2023b) detects poisoned samples in the training set by introducing confusing batches of benign data with randomly modified labels. The confusing batches with random labeling corrupt the benign correlations between normal semantic features and semantic labels, so the inference model trained with confusing batches and the poisoned dataset will find it hard to distinguish benign samples. However, the correlation between the backdoor trigger and the target label remains intact, as the confusing batches contain no trigger. Therefore, samples with correctly predicted labels by the inference model are considered poisoned. PDB (Wei et al., 2024) proactively injects a defensive backdoor into the model during training, overriding the potential backdoor injected by the poisoned training data. In summary, proactive defenses assume a stronger defender for better defensive performance.
@@
Figure 2. TAC (Zheng et al., 2022) analysis of different backdoor attacks on 512 neurons. The $y$ axis contains the TAC values, and the $x$ axis depicts the index of neurons. Higher TAC values suggest a stronger relation between corresponding neurons and the backdoor trigger. In the Appendix App. C.7, we also provide sorted TAC value plots, clearly showing the prominent TAC values.

### v3

@@ Abstract.
Model Inspection NC (wang_2019_nc)
Tabor (guo_2020_tabor)
FeatureRE (wang2022featurere)
Unicorn (wang2023unicorn)
BTI-DBF (xu2024btidbf)
Input Inspection Scale-up (guo2023scaleup)
IBD-PSC (hou2024ibdpsc)
CT (qi2023proactivedetection)
Pruning FP (liu2018finepruning)
ANP (Wu_2021_anp)
CLP (zheng2022clp)
RNP (li2023rnp)
@@ 1. Introduction
While deep neural networks (DNNs) have achieved excellent performance on various tasks, they are vulnerable to backdoor attacks. Backdoor attacks insert a secret functionality into a model, which is activated by malicious inputs during inference. Such inputs contain an attacker-chosen property called the trigger. Backdoored DNNs can be created by training with poisoned data (Gu_2019_badnet; chen_2017_blend; nguyen2021wanet; turner2019lc). More powerful and stealthy backdoors can also be injected through the control of a training process (shokri2020bypassing; Bagdasaryan_2021_blind; Cheng2021dfst; nan2022iba; Qi2022deploymentbackdoor; mo2024ssdt; zhao2022defeat), or by direct weights modification of the victim model (hong2022handcrafted; cao2024dfba).
@@
In early backdoor attacks (Gu_2019_badnet; chen_2017_blend; liu_2018_Trojannn), triggers could induce noticeable changes that human inspectors or anomaly detectors (chen2018detecting; wang_2019_nc; liu_2019_abs) could easily spot. To enhance the ability to remain undetected against such defenses (i.e., achieve input-space stealthiness), smaller or more semantic-aware triggers are designed (nguyen2021wanet; Doan2021lira; Wang_2022_bpp). Input-space stealthy backdoor attacks usually need to change labels of poisoned samples to the target class (i.e., dirty-label), which makes detection easier (chen2018detecting). To this end, another line of backdoor attacks poisons the training data without changing the labels (turner2019lc; zeng2023narcisus) (i.e., clean-label), improving backdoor stealthiness.
@@
Despite the stealthiness concerning input images and labels, it has been widely observed that existing backdoor attacks introduce separable representations in the feature space, which can be exploited to develop backdoor defenses (wang2022featurere; qi2022revisitingadaptive; min2023fst; Zhu2023ftsam; xu2024btidbf). For example, featureRE (wang2022featurere) utilizes feature separability and designs a feature space constraint to reverse-engineer the backdoor trigger. In response to feature-space defenses, state-of-the-art (SOTA) backdoor attacks focus on eliminating the separability in the feature space (mo2024ssdt; qi2022revisitingadaptive; abad2024gradientshaping; shokri2020bypassing) to increase the feature-space stealthiness, i.e., the undetectability against feature-space defenses. Considering a different threat model, supply-chain backdoor attacks assume control over the training or directly modify the model’s weights (Cheng2021dfst; hong2022handcrafted; cao2024dfba), and the backdoored model is provided as a service or as the final product. For example, supply-chain attacks could introduce a penalty to the training loss that decreases the distance between the backdoor and benign features to increase feature-space stealthiness (shokri2020bypassing; zhao2022defeat; doan2021wb; nan2022iba).
@@
An important observation is that most backdoor attacks are designed to be stealthy to resist a specific type of defense. For example, WaNet (nguyen2021wanet) and Bpp (Wang_2022_bpp) design imperceptible triggers to bypass input-space defenses (such as NC (wang_2019_nc)), but introduce significant separability in the feature space (wang2022featurere). Adap-patch (qi2022revisitingadaptive) avoids feature separability but uses patch-based triggers, which a human inspector can detect. More critically, current backdoor attacks are barely evaluated against parameter-space defenses (Zhu2023ftsam; lin2024unveiling; xu2024ban; Wu_2021_anp; zheng2022clp; li2023rnp). This oversight is significant because backdoor behaviors are ultimately embedded in and reflected by the parameters of the backdoored model, which is the final product of any backdoor attack. As such, there is a lack of systematic evaluation of backdoor attacks against the latest parameter-space defenses.
@@
Figure 1. Diagram illustrating the working mechanism of Grond. On the left, universal PGD (UPGD) perturbation is generated as backdoor patterns to be injected. In the middle, ABI is applied where perturbed samples are iteratively used to train the model, and the model parameters are pruned to limit the magnitude of prominent backdoored weights. On the right, the output backdoored model that considers comprehensive stealthiness is deployed, where 1) the triggers are invisible, 2) the features of trigger samples are inseparable, and 3) the backdoored model weights are hardly distinguishable from benign model weights. Perturbations generated by UPGD are scaled up $10\times$ for visualization.
@@ 2.1. Preliminaries on Backdoor Training
[…]
@@ 2.2. Backdoor Attacks
Backdoor attacks compromise the integrity of the victim model so that the model performs naturally on benign inputs but is misled to the target class by inputs containing the backdoor trigger. The trigger can be a visible pattern inserted into the model’s input in the input space or a property that affects the feature representation of the model’s input in the feature space. Eventually, however, the backdoored model’s parameters in the parameter space will be altered regardless of the exact backdoor attack (see Figure 2). To insert a backdoor, the attacker is assumed to only control a small portion of the training data under the poison training scenario (Gu_2019_badnet; chen_2017_blend; zhang2021advdoor). In the supply-chain setting (backdoor models provided to users), the attacker also controls the training process (shokri2020bypassing; nguyen2020inputaware; Bagdasaryan_2021_blind; nguyen2021wanet; Wang_2022_bpp). Moreover, the backdoor can also be created by directly modifying the model’s weights (liu2017faultinjection; hong2022handcrafted; Qi2022deploymentbackdoor; cao2024dfba).
@@
Input-space attacks. Traditional attacks typically use simple patterns as their triggers. For example, BadNets (Gu_2019_badnet) uses a fixed patch, and Blend (chen_2017_blend) mixes a Hello Kitty pattern into the images as the trigger. These non-stealthy triggers introduce abnormal data into training data and can be easily detected by human inspectors or defenses (chen2018detecting; wang_2019_nc). To improve the stealthiness, various triggers are proposed to achieve invisibility in the input space. IAD (nguyen2020inputaware) designed a dynamic solution in which the triggers vary among different inputs. WaNet (nguyen2021wanet) proposed the warping-based trigger, which is invisible to human inspection. Although these methods successfully build invisible triggers and bypass traditional defenses (wang_2019_nc), they still introduce separable features and can be detected by feature-space defenses (wang2022featurere; xu2024btidbf). These input-invisible attacks can be even more noticeable than input-visible attacks (BadNet, Blend) in the feature space (xu2024ban). We conjecture this is because they have fewer modifications on input pixels than input-visible attacks. Therefore, input-invisible attacks require more influential features to achieve a successful attack.
@@
Feature-space attacks. Knowing the vulnerability of input-space attacks against feature-space defenses, backdoor attacks are improved for feature-space stealthiness. A common threat model of this attack type is to assume additional control over the training process. For example, (shokri2020bypassing; doan2021wb; zhao2022defeat; nan2022iba) directly designed loss functions to minimize the difference between the backdoor and benign features. Aside from design loss penalties, TACT (di2021tact) and SSDT (mo2024ssdt) point out that source-specific (poison only the specified source classes) attack helps to obscure the difference in features between benign and backdoor samples. In addition, (qi2022revisitingadaptive) proposed Adap-blend and Adap-patch, which obscures benign and backdoor features by 1) including poisoned samples with the correct label, 2) asymmetric triggers (using a stronger trigger at inference time), and 3) trigger diversification (using diverse variants of the trigger during training). Unfortunately, existing attacks lack systematic evaluation against the latest defenses. For example, Adap-blend can be thoroughly mitigated by recent works (Zhu2023ftsam; xu2024btidbf; xu2024ban). In summary, feature-space attacks usually introduce visible triggers and cannot defeat the latest defenses.
@@
Supply-chain attacks. Supply-chain attacks are getting more attention due to their potential in real-world applications where backdoored models are provided as the final product to users. In supply-chain attacks, adversaries could control both training data and the training process. Note that feature-space attacks (shokri2020bypassing; liu2020compositeattack; doan2021wb; zhao2022defeat; ren2021Simtrojan; Cheng2021dfst; nan2022iba; xia2023mmdregularization; mo2024ssdt) with the assumption of control over the training process are a subset of supply-chain attacks, as their output is the backdoor model. In addition to training control, another kind of supply-chain attack directly adjusts the model’s weights in parameter space to introduce a backdoor, i.e., parameter-space attack. T-BFA (rakin2022tbfa), TBT (Rakin_2020_TBT), and ProFlip (chen2021proflip) explore modifying a sequence of susceptible bits of DNN parameters stored in the main memory (e.g., DRAM) to inject the backdoor. SRA (Qi2022deploymentbackdoor) and handcrafted Backdoor (hong2022handcrafted) directly modify a subset of models’ parameters to increase the logits of the target class. However, these attacks require a local benign dataset to guide the search for the subset of parameters to be modified. Data-free backdoor (lv2023datafree) releases the requirement of benign data by collecting substitute data irrelevant to the main task and fine-tuning using the substitute data. DFBA (cao2024dfba) further proposes a retraining-free and data-free backdoor attack by injecting a backdoor path (a single neuron from each layer except the output layer) into the victim model. In summary, supply-chain attacks focus on increasing the backdoor’s effectiveness without comprehensively considering parameter-space defenses.
@@ 2.3. Backdoor Defenses
Backdoor defenses can be classified into detection and mitigation. Detection refers to determining whether a model is backdoored (model detection(wang_2019_nc; liu_2019_abs; zhao2022defeat; wang2023unicorn; xu2024btidbf) or a given input is applied with a trigger (input detection(gao2019strip; guo2023scaleup; mo2024ssdt). Mitigation refers to erasing the backdoor effect from the victim model by pruning the backdoor-related neurons (pruning-based defenses) (liu2018finepruning; Wu_2021_anp; zheng2022clp; li2023rnp) or unlearning the backdoor trigger (fine-tuning-based defenses) (Zhu2023ftsam; zeng2022ibau; min2023fst; xu2024btidbf). In addition, recent works (li2021abl; qi2023proactivedetection; wei2024pdb) also consider the home-field advantage [^1] The defender has full control of the system and could access the training process. to design more powerful proactive defenses.
@@
Backdoor detection. Backdoor trigger reverse engineering (also known as trigger inversion) is considered one of the most practical defenses for backdoor detection as it can be applied to both poisoning training and supply-chain scenarios (wang2022featurere; wang2023unicorn; xu2024btidbf; xu2024ban), i.e., it is a post-training method. Specifically, trigger inversion works by searching for a potential backdoor trigger for a specific model. The model is determined as backdoored if a trigger is found, and the trigger can be used to unlearn the backdoor. The searching is implemented as an optimization process corresponding to the model and a local benign dataset. For example, NC (wang_2019_nc) firstly proposes trigger inversion for detection by optimizing the mask and pattern in the input space that can mislead the victim to the target class. This optimization is repeated for all classes. The model is considered backdoored if an outlier significantly smaller than the triggers for all other classes exists. Although methods similar to NC perform well against fixed patch trigger attacks, such as BadNets (Gu_2019_badnet) and Blend (chen_2017_blend), they may not be effective against input-stealthy attacks like WaNet (nguyen2021wanet). To address this problem, FeatureRE (wang2022featurere) moves trigger inversion from input space to feature space. Unicorn (wang2023unicorn) further proposes a transformation function for attacks in other spaces, such as numerical space (Wang_2022_bpp). Recent works (xu2024btidbf; xu2024ban) focus on exploring new optimization objectives that address the inefficiency problem of previous trigger inversion methods due to optimization over all classes. BTI-DBF (xu2024btidbf) trains a trigger generator by maximizing the backdoor feature difference between benign samples and their generated version (by the trigger generator) and minimizing the benign feature difference. BAN (xu2024ban) optimizes the noise on neuron weights rather than input pixels to activate the potential backdoor, which further improves both effectiveness and efficiency.
@@
Backdoor mitigation. Backdoor mitigation consists of fine-tuning and pruning, which are effective and do not assume knowledge of backdoor triggers. Pruning methods aim to find and remove backdoor-related neurons. FP (liu2018finepruning) eliminates dormant neurons on benign inputs and then fine-tunes the pruned network. ANP (Wu_2021_anp) searches for backdoor-related neurons by adding adversarial noise to neuron weights to activate the backdoor. RNP (li2023rnp) uses an unlearning and recovering process on benign data to expose backdoor neurons, as the recovering will force the backdoor neurons to be silent for the main benign task. Unlike these pruning methods guided by benign data, CLP (zheng2022clp) directly analyzes the Channel Lipschitzness Constant of the network and prunes the high Lipschitz constant channels in a data-free manner.
@@
Traditional fine-tuning as a defense usually needs trigger inversion methods to recover the trigger and then unlearn the trigger. For example, BTI-DBF(U) (xu2024btidbf) fine-tunes backdoor models using triggers recovered by their inversion algorithm. However, there is no guarantee that the recovered trigger is the true trigger for the backdoor. Recent works also consider fine-tuning without the trigger information but with prior human knowledge. For example, FT-SAM (Zhu2023ftsam) observes a positive correlation between the weight norm of neurons and backdoor-related neurons. Then, they propose a fine-tuning method to revise the large outliers of weight norms using Sharpness-Aware Minimization (SAM). I-BAU (zeng2022ibau) forms a min-max fine-tuning similar to adversarial training, where the inner maximizing searches for perturbations that mislead the model, and the outer minimizing is to keep the model’s capability on benign data. FST (min2023fst) assumes the backdoor and benign features should be disentangled and actively shifting features while fine-tuning by encouraging the discrepancy between the original backdoor model and the fine-tuned model.
@@
Proactive defense. Several methods have been proposed to exploit the home-field advantage, i.e., a stronger defender, for better defensive performance. ABL (li2021abl) proposes two techniques to avoid learning the backdoor task while training on the poisoned data: 1) trapping the loss value of each example around a certain threshold because backdoor tasks are learned much faster than the main task, and their loss decreases much faster. The samples with lower loss are recorded as poisoned samples; 2) unlearning the backdoor with the recorded poisoned samples. CT (qi2023proactivedetection) detects poisoned samples in the training set by introducing confusing batches of benign data with randomly modified labels. The confusing batches with random labeling corrupt the benign correlations between normal semantic features and semantic labels, so the inference model trained with confusing batches and the poisoned dataset will find it hard to distinguish benign samples. However, the correlation between the backdoor trigger and the target label remains intact, as the confusing batches contain no trigger. Therefore, samples with correctly predicted labels by the inference model are considered poisoned. PDB (wei2024pdb) proactively injects a defensive backdoor into the model during training, overriding the potential backdoor injected by the poisoned training data. In summary, proactive defenses assume a stronger defender for better defensive performance.
@@
Figure 2. TAC (zheng2022clp) analysis of different backdoor attacks on 512 neurons. The $y$ axis contains the TAC values, and the $x$ axis depicts the index of neurons. Higher TAC values suggest a stronger relation between corresponding neurons and the backdoor trigger. In the Appendix App. C.7, we also provide sorted TAC value plots, clearly showing the prominent TAC values.